OpenStudy (anonymous):

use the sum or difference identities to find the exact value of the cosine 105

OpenStudy (raden):

cos105 = cos(45+60) use the formula : cos(a+b) = cos(a)cos(b) - sin(a)sin(b) so, cos(45+60) = cos(45)cos(60) - sin(45)sin(60) = ... continue it